OpenStudy (kiki99):

Solve this equation, step by step, using the method mentioned below : 3x^2-12x-15=0 graph, such are: vertex, direction of opening, axis of symmetry, y-intercept, x-intercept(s)):

OpenStudy (kiki99):

@sleepyjess

OpenStudy (kiki99):

@Mehek14

OpenStudy (kiki99):

@Mahoganie.Carson

OpenStudy (kiki99):

@Er.Mohd.AMIR

OpenStudy (er.mohd.amir):

make it perfect squre

OpenStudy (er.mohd.amir):

x*x-4x-5=0

OpenStudy (er.mohd.amir):

(x*x-2*2*x+2*2)-2*2-5=0 (x-2)^2-4-5=0

OpenStudy (er.mohd.amir):

(x-2)^2=9

OpenStudy (kiki99):

where do you get 4, and 5 from

OpenStudy (er.mohd.amir):

since i add 2*2 so subtract it and 5 is already there

OpenStudy (er.mohd.amir):

i thick u write wrong Q here please check it again

OpenStudy (kiki99):

no its correct

OpenStudy (kiki99):

the 2 is an exponent

OpenStudy (er.mohd.amir):

there is must be any "y"

OpenStudy (kiki99):

im so confused. im going to call my teacher

OpenStudy (er.mohd.amir):

ok

OpenStudy (mahoganie.carson):

Sorry I forgot how to do this @kiki99 :(